# Adversarial Evasion Attacks Practicality in Networks: Testing the Impact of Dynamic Learning Authors’ draft for soliciting feedback: April 4, 2024

### v3

@@ Abstract
Machine Learning (ML) has become ubiquitous, and its deployment in Network Intrusion Detection Systems (NIDS) is inevitable due to its automated nature and high accuracy compared to traditional models in processing and classifying large volumes of data. However, ML has been found to have several flaws, most importantly, adversarial attacks, which aim to trick ML models into producing faulty predictions. While most adversarial attack research focuses on computer vision datasets, recent studies have explored the suitability of these attacks against ML-based network security entities, especially NIDS, due to the wide difference between different domains regarding the generation of adversarial attacks.
To further explore the practicality of adversarial attacks against ML-based NIDS in-depth, this paper presents three distinct contributions: identifying numerous practicality issues for evasion adversarial attacks on ML-NIDS using an attack tree threat model, introducing a taxonomy of practicality issues associated with adversarial attacks against ML-based NIDS, and investigating how the dynamicity of some real-world ML models affects adversarial attacks against NIDS. Our experiments indicate that continuous re-training, even without adversarial training, can reduce the effectiveness of adversarial attacks. While adversarial attacks can compromise ML-based NIDSs, our aim is to highlight the significant gap between research and real-world practicality in this domain, warranting attention.
Index Terms:
@@
machine learning, dynamic machine learning, adversarial attacks, network intrusion detection systems
I Introduction
@@
Machine Learning (ML) applications are present in every corner. ML found its way to network security applications due to its (1) automated nature and (2) ability to process and classify high volumes of data with high accuracy. Thus, ML deployment in Network Intrusion Detection Systems (NIDS) is a foregone conclusion. However, in recent years, researchers highlighted that ML has a lot of drawbacks and flaws. To name a few, ML has privacy concerns [1] and suffers from reproducibility crises (when the results of a model can not be replicated using the exact code and dataset) [2]. Nevertheless, one ML challenge shined the most in the research community; adversarial attacks. In a nutshell, these adversarial attacks aim to trick ML models, making them produce faulty predictions. There is no doubt that these attacks are a menace to ML. However, it was observed that the majority of adversarial attack research uses computer vision datasets to evaluate its severity [3]. Thus, recently, a few researchers [4] [5] [6] have explored the practicality of adversarial attacks against ML-based network security entities, especially Network Intrusion Detection Systems (NIDS), and implied that attacking them is much more complicated than computer vision.
@@
Contributions: In this paper, we study the practicality of evasion adversarial attacks against NIDS and make three distinct contributions; First, we identify the questionable practicality prerequisites required to carry out evasion adversarial attacks against NIDS. We introduce an attack tree, a threat modeling technique, to visualize various scenarios attackers might use to launch evasion adversarial attacks against ML-NIDS. The attack tree helps identify leaf nodes with questionable feasibility. Secondly, we extend the research discussion on the practicality of adversarial attacks against NIDS by introducing a taxonomy of practicality issues associated with adversarial attacks against ML-based NIDS. This taxonomy includes issues reported in the literature, along with our own questions and observations. The third contribution emerged from our observation that the overwhelming bulk of adversarial attack research deals with ML approaches as static models. However, nowadays, numerous ML models are dynamic and constantly evolving. Approaches like stream learning [7] or online machine learning [8] constantly update their models to make them adaptive to new cases and data (prevent concept drifts). Therefore, we explore and test the effect of continuous training on adversarial attacks against NIDS. To the best of our knowledge, these are the first experiments on the impact of continuous training on such invasions against ML-based NIDS. This paper builds upon the preliminary findings presented in our poster [6].
The rest of the paper is organized as follows: Section II delves into the relevant background. In Section III, we review previous research and surveys. The attack tree, which shows the various techniques attackers might employ to launch evasion-based adversarial attacks against ML-NIDS, is presented in Section IV. Section V introduces the taxonomy of practical challenges and issues related to adversarial attacks against ML-NIDS. Our investigation environment and methodology are presented in Section VI. The outcomes of our study, illustrating the impact of continuous training on adversarial attacks in the NIDS domain, are outlined in Section VII. Section VIII encompasses discussions and lessons learned from our research. Finally, Section IX concludes our work and provides insights into future directions.
II Background
@@
II-A ML-based NIDS
II-B Dynamic Machine Learning
Concept drift in machine learning is the unforeseeable statistical changes in the prediction target’s properties [12]. In other words, the norm changes over time, so the ML model makes wrong predictions. In an adversary environment, where attackers change their techniques all the time, like the network/cybersecurity domain, concept drifts will more than likely occur [7]. Thus, to counter these drifts ML models must be dynamic, i.e, continuously evolving. Numerous terms are used to describe dynamic ML models (e.g., Online ML, Evolving ML, Continuous/Continual ML, and Stream ML). In the ML community, the implementation and definition of these kinds of ML techniques could differ from one study to another. However, the following concept is common; "The initial training set loses its validity as time passes due to changes of conditions in the aimed task. Thus, a mechanism to update a given model in order to adapt to new conditions is necessary" [8]. In simple terms, dynamic ML models must be updated and re-trained all the time to resist the effect of concept drift. There are different techniques for re-training (or continuous training); for example, adaptive windowing, where the model is re-trained when a concept drift is detected [13]. However, in practice, there are more elementary solutions, with less overhead on the system, like periodically re-training the model with recent data. In our tests, we use the latter approach.
II-C Adversarial Attacks Against Network Security and Defenses
@@
II-C1 Taxonomy of Adversarial Attacks in Network Security
Attackers’ Knowledge:
In White-box Attacks, the attackers have complete knowledge of the attacked ML model (which ML algorithm was used, gradient, hyperparameters, etc.). The complete opposite is Black-box Attacks, where the attackers have no knowledge of the attacked ML model. However, in most cases, researchers assume that the attackers crafting black-box attacks have access to the target ML system’s output. Thus, the attackers can create a substitute model (pretty close to the attacked model) by querying the targeted system. Then they use that substitute system to get the knowledge they need to attack. Grey-box Attacks lie between them (the attackers have "some" knowledge of the attacked ML system).
Attackers’ Approaches:
In Evasion Attacks, which is the main focus of this paper, attackers aim to deceive the ML model during the decision-making process by adding carefully crafted perturbations to the input data. Like [16], where a perturbation is added to the features to make the IoT ML intrusion detection model output the wrong classification. As for Poisoning Attacks, the attackers add malicious samples to the training datasets (dataset contamination). Like [17], where a data poisoning attack against a federated learning-based IoT intrusion detection system allowed the adversaries to implant a backdoor into the ML detection model to wrongly classify malicious traffic as benign. In Model Stealing Attacks, an attacker can probe a black box model to extract all the vital information to create a "clone" model very close to the target model. It is worth mentioning that ML models in a monumental system like NIDS are very expensive, and requires a lot of resources to create. Thus, stealing them will be a massive loss to the associations that created them. The main focus of this paper is evasion attacks. Regarding Backdoor Attacks, the attackers aim to embed a backdoor into the ML model. The backdoored model operates normally when processing inputs lacking the trigger. However, it is misdirected to perform the attacker’s sub-task once the secret trigger is presented in the input [18].
@@ II-C1 Taxonomy of Adversarial Attacks in Network Security
Okunle Ibitoye et al. [19] presented the space term in adversarial attacks taxonomy in the network security domain. In Feature-space attacks, the attacker only modifies (by adding perturbations) the ML model input feature vector to craft the attack. As for the Problem-space attacks, the attacker modifies the actual file (malware, pcap, PE (Portable Executable), etc.) to trick the attacked model.
II-C2 Defenses Against Adversarial Attacks
There are multiple approaches to defending ML models against adversarial attacks. Here are a couple of examples: Adversarial Training: This involves inserting adversarial examples into the training dataset to enhance the model’s robustness against such attacks [20] [21]. Feature Reduction: Reducing the feature sets may decrease the effectiveness of perturbations, thereby making the ML classifier less vulnerable to evasion attacks. This technique has shown promising results in ML spam and malware detection models [22].

### v5

# Evasion Adversarial Attacks Remain Impractical Against ML-based Network Intrusion Detection Systems, Especially Dynamic Ones

@@ Abstract
Machine Learning (ML) has become pervasive, and its deployment in Network Intrusion Detection Systems (NIDS) is inevitable due to its automated nature and high accuracy compared to traditional models in processing and classifying large volumes of data. However, ML has been found to have several flaws, most importantly, adversarial attacks, which aim to trick ML models into producing faulty predictions. While most adversarial attack research focuses on computer vision datasets, recent studies have explored the suitability of these attacks against ML-based network security entities, especially NIDS, due to the wide difference between different domains regarding the generation of adversarial attacks. To further explore the practicality of adversarial attacks against ML-based NIDS in-depth, this paper presents several key contributions: identifying numerous practicality issues for evasion adversarial attacks on ML-NIDS using an attack tree threat model, introducing a taxonomy of practicality issues associated with adversarial attacks against ML-based NIDS, identifying specific leaf nodes in our attack tree that demonstrate some practicality for real-world implementation and conducting a comprehensive review and exploration of these potentially viable attack approaches, and investigating how the dynamicity of real-world ML models affects evasion adversarial attacks against NIDS. Our experiments indicate that continuous re-training, even without adversarial training, can reduce the effectiveness of adversarial attacks. While adversarial attacks can compromise ML-based NIDSs, our aim is to highlight the significant gap between research and real-world practicality in this domain, which warrants attention.
@@
Keywords: Machine learning; Dynamic machine learning; Adversarial attacks; Network intrusion detection systems
1 Introduction
@@
Machine Learning (ML) applications are present in every corner. ML found its way to network security applications due to its (1) automated nature and (2) ability to process and classify high volumes of data with high accuracy. Thus, ML deployment in Network Intrusion Detection Systems (NIDS) is a foregone conclusion. However, in recent years, researchers highlighted that ML has a lot of drawbacks and flaws [1] [2], nevertheless, one ML challenge stood out the most in the research community; adversarial attacks. In a nutshell, these adversarial attacks aim to trick ML models, making them produce faulty predictions. There is no doubt that these attacks are a menace to ML. However, it was observed that the majority of adversarial attack research uses computer vision datasets to evaluate its severity [3]. Thus, recently, a few researchers [4] [5] have explored the practicality of adversarial attacks against ML-based network security entities, especially Network Intrusion Detection Systems (NIDS), and implied that attacking them is much more complicated than computer vision.
@@
Contributions: In this paper, we study, evaluate, and contextualize the practicality of evasion adversarial attacks against NIDS and make several key contributions. First, we identify the questionable practicality prerequisites required to carry out evasion adversarial attacks against NIDS. We introduce an attack tree, a threat modeling technique, to visualize various scenarios attackers might use to launch evasion adversarial attacks against ML-NIDS. The attack tree helps identify leaf nodes with questionable feasibility. Second, we extend the research discussion on the practicality of adversarial attacks against NIDS by introducing a taxonomy of practicality issues associated with adversarial attacks against ML-based NIDS. This taxonomy includes issues reported in the literature, along with our own questions and observations. Third, we identify specific leaf nodes in our attack tree that demonstrate some practicality for real-world implementation and conduct a comprehensive review and exploration of these potentially viable attack approaches. Fourth, our contribution emerged from our observation that the overwhelming bulk of adversarial attack research deals with ML approaches as static models. However, nowadays, numerous ML models are dynamic and constantly evolving. Approaches like stream learning [6] or online machine learning [7] constantly update their models to make them adaptive to new cases and data (prevent concept drifts). Therefore, we explore and test the effect of continuous training on gradient-based adversarial attacks against NIDS. To the best of our knowledge, these are the first experiments on the impact of continuous training on such invasions against ML-based NIDS.
Table 1: Comparative Analysis of our paper and Related Work on Adversarial Attacks in Network Intrusion Detection Systems ($\bullet$ = full, = partial, $\circ$ = no coverage).
Paper Evasion Adversarial Attacks on NIDS Threat Modeling (Attack Trees) & Feasibility Assessment Practicality Analysis of Attacks Taxonomy of Attack Practicality Issues Impact of Dynamic Learning
[…]
Note: Feasibility Assessment includes identifying attack tree leaves with questionable and possible feasibilities
2 Related Work
2.1 Adversarial attacks against ML-NIDS
Numerous researchers have examined adversarial attacks against machine learning-based network intrusion detection systems. Alatwi et al. [10] surveyed and categorized the studies of adversarial attacks against ML-based NIDSs and briefly discussed their applicability in real-world scenarios. He et al. [9] has also conducted a comprehensive exploration of white- and black-box adversarial attacks on DNNs, and their potential relevance in NIDS. They also analyzed existing defense mechanisms against these attacks. Furthermore, Martins et al. [12] examined adversarial machine learning against intrusion and malware detection systems. They concluded that a wide variety of the reviewed attacks could be effective, however, their practicality was not been tested enough in intrusion scenarios. Furthermore, They have stated that some of the datasets in intrusion scenarios are outdated, and most of them are similar and lack variety. Ennaji et al. [8] also surveyed evasion adversarial attacks on NIDS and provided practical recommendations for adapting adversarial strategies to network-specific constraints.
2.2 Practicality of Adversarial attacks against ML-NIDS
Some studies have focused on the practicality of adversarial attacks in the network intrusion detection domain; Vitorino et al. [11] presented an overview of adversarial learning techniques capable of generating realistic examples to deceive NIDSs. Merzouk et al. [5] also attempted to inspect the practicality of adversarial attacks to evade network intrusion detection systems. They highlighted some limitations of adversarial attacks against ML-NIDS. However, they have only concentrated on the feature space attacks. Apruzzese et al. [4] aimed to specify and model the abilities and circumstances necessary for attackers to execute practical adversarial attacks on network intrusion detection systems. Both [11] and [4] primarily reviewed and evaluated attacks from research literature but did not consider the dynamic nature of modern ML models. In a separate study, Apruzzese et al. [3] emphasized the existence of a gap between researchers and practitioners in the field of adversarial attacks. Their primary focus was not on NIDSs; however, through collaboration between industry and academia, they articulated the general viewpoint that “Real attackers don’t compute gradients.” In other words, they were unable to identify an attacker who computes gradients to craft real-world adversarial attacks.
2.3 Comparison with Related Work
As seen in Table 1, which shows the comparative analysis of our paper and related work on adversarial attacks in the NIDS domain, our work, unlike other relevant related work, covers all the comparison points in the list. To the best of our knowledge, none of the previous work has explored and tested the practicality of adversarial attacks against dynamic ML models in the intrusion detection domain. Moreover, to the best of our knowledge, none of the previous work has merged systematic literature review, threat modeling techniques, and testing with evasion adversarial attacks on NIDS to evaluate the practicality of evasion adversarial attacks in the NIDS domain. Alatwi et al. [13] conducted threat modeling techniques for adversarial attacks against NIDS; however, their attack tree covered all adversarial attacks, while ours focuses on just evasion attacks, so naturally ours was more in-depth, covered more angles, and identified attack tree leaves with questionable and possible feasibilities. Unlike their work, ours was backed by systematic literature review and testing.
3 Methodology of Systematic Literature Review
We conduct a systematic literature review following a six-phase approach as shown in Figure 1. Our methodology begins by defining the scope to focus on evasion attacks affecting ML/DL NIDS published in 2022 or after. We design a systematic search query using three search terms in conjunction with a temporal constraint (Table 2): adversarial attack terms, network intrusion detection terms, and machine learning terms. The search is conducted across IEEE Xplore, ACM Digital Library, ScienceDirect/Elsevier, ArXiv, and Google Scholar. Following our initial title and abstract screening using the criteria in Table 3, we conduct a second, thorough screening to remove duplicates, repeated contributions, and articles with limited novelty. We then perform full-text screening and extract key data including threat models, knowledge requirements, and implementation constraints. Finally, we use the gathered data to construct an attack tree, identify practicality issues in current attacks, define requirements for practical adversarial attacks on NIDS, and systematically classify papers that meet our practicality criteria.
The temporal constraint of focusing on publications from 2022 onwards is motivated by He et al.’s [9] comprehensive survey in IEEE Communications Surveys & Tutorials, which stated that “existing adversarial attacks lack a comprehensive evaluation of the maliciousness of the adversarial examples” and highlighted insufficient practical implementation considerations in adversarial attacks against NIDS. This critical assessment of pre-2022 research motivates our investigation: we aim to determine whether adversarial attacks published from 2022 onwards have become more practical and realistic in addressing real-world deployment constraints, or if these fundamental limitations persist in recent literature.
Figure 1: Systematic Literature Review
Table 2: Search query format used to retrieve papers. The query is composed of three search terms in conjunction with a temporal constraint. The values in the second column replace each search term.
Search Query <search_term_1> AND <search_term_2> AND <search_term_3> AND <temporal_constraint>
<search_term_1> “adversarial attack” OR “adversarial example” OR “adversarial perturbation” OR “adversarial machine learning” AND “evasion”
<search_term_2> “network intrusion detection” OR “intrusion detection system” OR “NIDS” OR “IDS”
<search_term_3> “machine learning” OR “ML”
<temporal_constraint> Publication year: 2022 or after
Table 3: Inclusion and Exclusion Criteria for Systematic Literature Review
Criteria Type Description
Inclusion Criteria • Published between 2022 and 2026
• Focuses on evasion adversarial attacks
• Targets ML/DL-based network intrusion detection systems
• Written in English
Exclusion Criteria • Targets defenses, not evasion adversarial attacks
• Focuses on poisoning attacks, model stealing, or backdoor attacks only
• Targets domains other than network intrusion detection
• Insufficient technical detail for analysis
• Duplicate publications (only most complete version retained)
4 Background
4.1 ML-based NIDS
4.2 Dynamic Machine Learning
Concept drift in machine learning is the unforeseeable statistical changes in the prediction target’s properties [17]. In other words, the norm changes over time, so the ML model makes wrong predictions. In an adversary environment, where attackers change their techniques all the time, like the network/cybersecurity domain, concept drifts will more than likely occur [6]. Thus, to counter these drifts ML models must be dynamic, i.e, continuously evolving. Numerous terms are used to describe dynamic ML models (e.g., Online ML, Evolving ML, Continuous/Continual ML, and Stream ML). In the ML community, the implementation and definition of these kinds of ML techniques could differ from one study to another. However, the following concept is common; “The initial training set loses its validity as time passes due to changes of conditions in the aimed task. Thus, a mechanism to update a given model in order to adapt to new conditions is necessary” [7]. In simple terms, dynamic ML models must be updated and re-trained all the time to resist the effect of concept drift. There are different techniques for re-training (or continuous training); for example, adaptive windowing, where the model is re-trained when a concept drift is detected [18]. However, in practice, there are more elementary solutions, with less overhead on the system, like periodically re-training the model with recent data. In our tests, we use the latter approach.
@@ 4.2 Dynamic Machine Learning
4.3 Adversarial Attacks Against Network Security and Defenses
@@
4.3.1 Taxonomy of Adversarial Attacks in Network Security
Attackers’ Approaches: In Evasion Attacks, which is the main focus of this paper, attackers aim to deceive the ML model during the decision-making process by adding carefully crafted perturbations to the input data. Like [21], where a perturbation is added to the features to make the IoT ML intrusion detection model output the wrong classification. As for Poisoning Attacks, the attackers add malicious samples to the training datasets (dataset contamination). Like [22], where a data poisoning attack against a federated learning-based IoT intrusion detection system allowed the adversaries to implant a backdoor into the ML detection model to wrongly classify malicious traffic as benign. In Model Stealing Attacks, an attacker can probe a black box model to extract all the vital information to create a “clone” model very close to the target model. It is worth mentioning that ML models in a monumental system like NIDS are very expensive, and requires a lot of resources to create. Thus, stealing them will be a massive loss to the associations that created them. The main focus of this paper is evasion attacks. Regarding Backdoor Attacks, the attackers aim to embed a backdoor into the ML model. The backdoored model operates normally when processing inputs lacking the trigger. However, it is misdirected to perform the attacker’s sub-task once the secret trigger is presented in the input [23].
Figure 3: Attackers’ Knowledge Types
Attackers’ Knowledge: As seen in Figure 3, there are three types of knowledge in adversarial attacks: white-, gray-, and black-box attacks. In White-box Attacks, the attackers have complete knowledge of the attacked ML model (which ML algorithm was used, gradient, hyperparameters, model architecture, parameters, etc.). The complete opposite is Black-box Attacks, where the attackers have no knowledge of the attacked ML model’s internals. There are query-based and non-query-based black-box attacks. The query-based attacks can be score-based, where the attacker queries the model to obtain the confidence score to get an approximation to the decision boundary of the target model, or decision-based, where attackers only have access to the final model decision (e.g., class label) without any confidence scores or probability information to create a surrogate model where they can then attack with white-box attack methods. Query-based attacks present significant practicality limitations as outsiders don’t have access to NIDS outputs, so they cannot get confidence scores nor decisions. Non-query-based black-box attacks, the most practical attacks, can be blind or random non-gradient perturbations to input network traffic, which produce inconsistent results, or non-query-based transfer-based attacks, where an attacker trains a surrogate model on the same data distribution as the target model to launch black-box attacks using the transferability property of ML models. Gray-box Attacks lie between white-box and black-box scenarios, where attackers have partial knowledge of the attacked ML system.
@@ 4.3.1 Taxonomy of Adversarial Attacks in Network Security
Figure 4: Feature-Space vs. Problem-Space Perturbations in NIDS Domain: If modifications occur at 1* to the input network flow, this constitutes a problem-space attack; if perturbations occur at 2* to the input feature vector, this constitutes a feature-space attack. In problem-space attacks, modifications such as increasing the size of one packet (the red one, for example) to alter the Max Packet Length feature must be mapped to the ML model’s input feature vector. This process is called inverse feature mapping [24]. Note that the perturbations must remain malicious after both feature extraction and pre-processing (red perturbations), not just after feature extraction alone (purple perturbations).
@@
Ibitoye et al. [25] presented the space term in adversarial attacks taxonomy in the network security domain. In Feature-space attacks, the attacker only modifies (by adding perturbations) the ML model input feature vector to craft the attack. As for the Problem-space attacks, the attacker modifies the actual file (malware, pcap, PE (Portable Executable), etc.) to trick the attacked model. Figure 4 shows the difference between feature-space and problem-space Perturbations in the NIDS domain.
4.3.2 Defenses Against Adversarial Attacks
There are multiple approaches to defending against adversarial attacks; as seen in Figure 5 we classify them into 3 classes: pre-processing techniques, ML-model hardening techniques, and post-processing techniques. (1) Pre-processing techniques include defense methods preceding ML model inference like Adversarial Detection; for example, Ren et al. [26] proposed adversarial attack detection methods by using the causal inference technique. He et al. [9] classified Adversarial Detection techniques into secondary classifier, projection-based, statistics-based, and mutation-based, which is the most practical because it is applicable and generalizable [9]. Feature Reduction might decrease the effectiveness of perturbations, thereby making the ML classifier less vulnerable to evasion attacks as it reduces the attack-space. Elshehaby et al. [27] took it further by applying domain-specific constraints rather than blindly or randomly reducing the feature vector. They used their novel score to select resilient features during feature selection or mask non-resilient features. Input Randomization involves adding random padding on inputs to lessen the effect of adversarial perturbations [25]. Adversarial Query Detection aims to defend against query-based black-box adversarial attacks by detecting these queries. Usually, these defenses analyze query patterns and sequences to identify suspicious behavior in queries received by the system [28]. (2) ML-model hardening techniques involve defense methods that make the ML model more robust against evasion adversarial attacks; these methods are implemented during model development or training like Adversarial Training which involves inserting adversarial examples into the training dataset to enhance the model’s robustness against such attacks [29] [30]. Gradient Masking modifies machine learning models in an attempt to obscure their gradient from an attacker [31] since most evasion adversarial attacks are based on the use of gradient. Gradient regularization is a technique that constrains gradient magnitudes across neural network layers to improve model robustness and prevent overfitting. This approach has been studied as a defense against adversarial attacks, with researchers like Ross et al. [32] demonstrating that gradient regularization enhances deep neural network resilience to adversarial perturbations. (3) Post-processing techniques involve defense methods post ML model inference and development. Output randomization is a defense strategy that exploits the limited knowledge available to black-box attackers, who can only access model confidence scores without knowing the underlying architecture. Kwon et al. [33] developed AdvGuard, which injects noise into the softmax layer to randomize confidence values returned to queries. This randomization disrupts black-box attacks that depend on consistent confidence feedback, making adversarial example generation infeasible by providing unreliable scoring information. Knowledge distillation, initially developed by Hinton et al. [34] for compressing large neural networks into smaller ones, was later repurposed by Papernot et al. [35] as an adversarial defense mechanism. This adapted approach uses the original network’s outputs as training targets for a distilled model to reduce overfitting and increase the robustness of the model. Similar to ensemble learning, researchers have proposed combining multiple defense strategies to enhance robustness against adversarial attacks (Ensemble Defenses). Adaptive Continuous Adversarial Training (ACAT) [36][37] exemplifies this approach by merging adversarial training with detection, incorporating newly detected adversarial examples into ongoing training processes.
Figure 5: Defenses Against Adversarial Attacks
While these defense mechanisms might provide protection against adversarial attacks, they suffer from significant limitations. The most critical drawback is the substantial system complexity and associated costs, requiring additional computational resources, specialized expertise, and ongoing maintenance overhead. Pre-processing techniques like adversarial detection and input randomization introduce latency, computational burden, and may degrade model performance on legitimate traffic. ML-model hardening approaches such as adversarial training dramatically increase training time and computational requirements, often reducing model accuracy. Post-processing defenses add another layer of computational complexity and may introduce inconsistent outputs affecting system reliability.
5 Identifying Infeasible and Feasible Adversarial Attack Prerequisites Through an Attack Tree Threat Model
The majority of the adversarial attacks were designed and tested in the literature using computer vision datasets [3]. Because of the focus on the computer vision datasets, it is not clear if these attacks will be applicable in the same way to the network/cybersecurity domain which has more adversaries and more security constraints and measures, making it much more complicated to evade. This contrast is depicted in Figure 6, which illustrates the crafting of evasion adversarial attacks in both the computer vision and network domains. Thus, in this section, we present an attack tree, depicted in Figure 7, that showcases the various strategies attackers might employ to execute evasion adversarial attacks. Additionally, we identify leaf nodes (prerequisites) with questionable feasibility marked by the ’?’ sign, and red cells denote leaf nodes with some feasibility. OR nodes represent disjunctions between alternative approaches, while conjunctions indicate required simultaneous conditions.